OpenStudy (anonymous):

Help will fan or medal! Evaluate 4(2-5)^3+1

Vocaloid (vocaloid):

start with the parentheses first (2-5) = -3, with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

now look at the exponent (-3)^3 = ?

OpenStudy (anonymous):

i got -27

Vocaloid (vocaloid):

good. now we multiply (-27)*4 = ?

OpenStudy (anonymous):

-108

Vocaloid (vocaloid):

good, now we just add 1 to that to get -107 and that's it

Vocaloid (vocaloid):

remember the order of operations: parentheses, exponents, multiplication/division, addition/subtraction

OpenStudy (anonymous):

@Vocaloid I got -107 as my final answer

Vocaloid (vocaloid):

good